# Wasserstein Distributionally Robust Nash Equilibrium Seeking with Heterogeneous Data: A Lagrangian Approach

### v1

@@ Abstract
We study a class of distributionally robust games where agents are allowed to heterogeneously choose their risk aversion with respect to distributional shifts of the uncertainty. In our formulation, heterogeneous Wasserstein ball constraints on each distribution are enforced through a penalty function leveraging a Lagrangian formulation. We then formulate the distributionally robust Nash equilibrium problem and show that under certain assumptions it is equivalent to a finite-dimensional variational inequality problem with a strongly monotone mapping. We then design an approximate Nash equilibrium seeking algorithm and prove convergence of the average regret to a quantity that diminishes with the number of iterations, thus learning the desired equilibrium up to an a priori specified accuracy. Numerical simulations corroborate our theoretical findings.
@@ 1.1 Motivation
Modern engineering applications are increasingly operated by autonomous agents that interact strategically to optimize interdependent objectives subject to constraints, thus giving rise to a game-theoretic setting. Recent developments in control, optimization, and machine learning rely heavily on a data-driven multi-agent framework. As some examples, power systems Sun et al. (2015); Molzahn et al. (2017) and networked robotics Bullo et al. (2009), Cortés and Egerstedt (2017), typically involve a high number of interacting decision-makers whose objectives are coupled through shared resources or performance constraints. At the same time, the environment these agents operate in is often uncertain, and its accurate probabilistic description is rarely known in practice. Thus, probabilistic models are often estimated from a finite collection of data, while statistical and algorithmic analyses are performed for the proposed methods. For uncertain multi-agent games Fabiani et al. (2022); Fele and Margellos (2021, 2019); Paccagnan and Campi (2019); Pantazis et al. (2024b) perform a detailed distribution-free probabilistic analysis for different game settings. From an algorithmic viewpoint, Franci and Grammatico (2021); Franci and Grammatico (2022b) design distributed schemes for stochastic generalized Nash equilibrium seeking. However, modeling uncertainty based on an empirical/ estimated distributions calculated on the basis of available samples and treated as ground truth can lead to lack of robustness towards even modest distribution shifts caused, for example, by changing operating conditions, sensor noise, seasonal demand variations or adversarial attacks. Thus, the decisions made are no longer (almost) optimal or as robust as expected.
@@
Distributionally robust optimization (DRO) was developed precisely to tackle this mismatch between the true distribution and its estimated counterpart. In particular, instead of optimizing for one nominal distribution, DRO optimizes for the worst-case distribution within a considered ambiguity set. This equips the decision with quantifiable out-of-sample and finite-sample performance guarantees. Thanks to duality and optimal transport tools Mohajerin Esfahani and Kuhn (2018); Aghassi and Bertsimas (2006) even when the ambiguity set is a Wasserstein ball centered at the empirical distribution, the resulting reformulation can be tractable, oftentimes formulated in the form of large linear programs under relatively mild assumptions. Classical DRO, however, is centralized, i.e., it assumes a single decision-maker and a single source of data. In contrast, in networked and learning-enabled systems, data are heterogeneous and private. Furthermore, as certain agents have different access to information/data, they might exhibit different levels of risk-aversion. A significant challenge that arises in several multi-agent systems is that agents can exhibit self-interested behaviour, wishing to minimize their own cost instead of a common social welfare cost. This framework gives rise to heterogeneous distributionally robust games, where each agent hedges against their own distributional ambiguity while interacting strategically with others.
@@
In Wang et al. (2022), the authors develop a no-regret learning algorithm for multi-agent convex games under CVaR-based risk measures. The work in Wang et al. (2025), studies convex games with heterogeneous feedback, while Wang et al. (2024) develops learning dynamics that converge to Nash equilibria in risk-averse games, where players optimize CVaR-type objectives. The work Pantazis et al. (2023) relied on epigraphic reformulation methods to provide tractable reformulations for such games. When the agents are affected by a common ambiguity set based on shared data and ambiguity radius, the results in Pantazis et al. (2023) recast the original DRNEP in a form which can provably be solved efficiently via centralized, decentralized or distributed algorithmic schemes. However, when ambiguity sets are private, this approach often leads to constraints which are private and coupling. An approach to fix this issue is to share the private constraints among agents. Unfortunately, this would introduce further cooperation among players, thus changing the original non-cooperative game with no clear connection between the equilibria of the two problems. To solve this problem, Pantazis et al. (2023) propose a primal-dual algorithm inspired by the golden-ratio method, Malitsky (2020), Franci and Grammatico (2022a), to empirically converge to an equilibrium of the original problem. Following a different approach, Pantazis et al. (2024a) focus on a particular class of DRNEPs and recast the problem into a variational inequality and propose efficient algorithmic methods to calculate an equilibrium. The recent paper in Alizadeh et al. (2025) follows again a VI formulation, but instead proposes a gradient descent/ascent for which regret guarantees can be obtained under certain assumptions. Lanzetti et al. (2025) introduce an alternative equilibrium concept based on optimal transport that protects against uncertainty in opponents’ behavior, while Fabiani and Franci (2023) study distributionally robust generalized Nash games over Wasserstein distributionally robust chance-constraints.
@@ 1.2 Main contributions
    Lagrangian formulation of DRNEPs with heterogeneous ambiguity sets: To improve the computational efficiency in DRNE equilibrium seeking, while still maintaining certain distributional robustness properties, we propose a Lagrangian formulation of DRNEPs, where each agent enforces its Wasserstein ambiguity constraint via a penalty term in their cost, thus extending the formulation of the single-agent setting proposed in Sinha et al. (2017). Our formulation enables heterogeneous and private ambiguity sets and risk-aversion levels without sharing constraints or data, while preserving the noncooperative nature of the game. We formalize the distributionally robust Nash equilibria (DRNE) for this novel setting and establish a connection with the DRNE of the original distributionally robust game. In our setting the agents’ penalty parameters act as individual distributional robustness-tuning coefficients.
@@ 2 Problem Formulation
Note that the inner supremum in Definition 2 involves selecting an adversarial probability distribution that maximizes the cost of each agent. The seemingly infinite dimensional problem $(G)$ can be shown to admit tractable reformulations through duality and the use of epigraphic reformulations in Pantazis et al. (2023) based on a methodology in the same spirit of Mohajerin Esfahani and Kuhn (2018). In particular, leveraging Kantorovich–Rubinstein duality, $(G)$ admits under mild assumptions the following reformulation:
@@
To circumvent this challenge, in this work we propose an alternative reformulation based on the extended Lagrangian method. Specifically, instead of optimizing the Lagrange multiplier of the Wasserstein constraint, we instead impose it as a penalty in each cost. As such, each agent $i\in\mathcal{N}$ solves the following optimization problem:
$\min_{x_{i}\in X_{i}}\max_{\mathbb{P}_{i}}\mathbb{E}_{\xi_{i}\sim\mathbb{P}_{i}}[f_{i}(x_{i},x_{-i},\xi_{i})]-\lambda_{i}W(\hat{\mathbb{P}}_{K_{i}},\mathbb{P}_{i})$ (6)
@@
Proof: The proof follows by adaptation of Sinha et al. (2017) for each agent $i\in\mathcal{N}$, given $x_{-i}$. $\blacksquare$
@@
For the game $G_{L}$ in (6), a point $x^{\ast}_{K}$ is distributionally robust Nash equilibrium (DRNE) if
$H_{i}(x_{K}^{\ast})$ $\leq\min_{x_{i}\in X_{i}}H_{i}(x_{i},x^{\ast}_{-i,K}),$ (8)
@@
Note that the subscript $K$ denotes that the decision depends on all the drawn samples.
@@ 3.1 Distributionally Robust Nash Equilbrium Problems as a Variational Inequality Problem
In this section we recast the game $G_{L}$ in (6) as a variational inequality and establish the conditions that guarantee (strong) monotonicity of the corresponding mapping. We propose an equilibrium seeking algorithm to calculate efficiently a DRNE of (7). In particular, we propose Algorithm 1 which is based on an inexact prox operation to calculate the inner supremum followed by a projected gradient descent for each agent $i\in\mathcal{N}$. We show that under certain conditions, convergence of the iterations to a neighbourhood of the equilibrium is guaranteed.
@@
Under Assumption 3.1a), $h_{i}^{(k_{i})}(x,\xi_{i})$ is $2\lambda_{i}$-strongly concave in $\xi_{i}$. $\square$
@@
Given Assumptions  3.1, we have $H_{i}$ is differentiable and satisfies
@@
Proof: The proof is an adaptation of Lemma 1 in Sinha et al. (2017) applied to the cost function of each agent $i\in\mathcal{N}$. $\blacksquare$
@@
Consider Assumptions 3.1 and 3.1. Then, the solution set of $VI(F_{K},X)$ is equivalent to the set of DRNE of $G_{L}$ in (6). $\square$
Proof: The proof can be obtained by applying Proposition 1.4.2 in Facchinei and Pang (2003) to VI($F_{K},X$). $\blacksquare$
@@
where $\xi=\operatorname*{col}((\xi)_{i=1}^{N})$. We then consider the following assumption:
@@
$G$ is $\mu$-strongly monotone in its first argument for any $\xi_{i}\in\Xi_{i}$, i.e, for any decisions $x,y\in X$ and $\xi_{i}\in\Xi_{i}$ it holds that
@@
$\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{i}H_{i}(x,\hat{\xi}_{i})-\nabla_{i}H_{i}(y,\hat{\xi}_{i}),x_{i}-y_{i}\Big\rangle$
$=\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{i}f_{i}(x,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{i}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i})),x_{i}-y_{i}\Big\rangle$
$=\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{i}f_{i}(x,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{i}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))$
$\quad+\nabla_{i}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{i}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i})),x_{i}-y_{i}\Big\rangle$
@@
$+\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{i}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{i}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i})),x_{i}-y_{i}\Big\rangle$
@@
$\left\|\nabla_{i}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{i}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i}))\right\|$
@@ 3.2 Average Convergence Guarantees
Let the assumptions above hold. Then, the iterates $(x_{t})_{t\in\mathbb{N}}$ of Algorithm 1 averagely converge to an $\epsilon$-DRNE, i.e.,
@@
$\left\|\nabla_{i}f_{i}(x_{i,t},x_{-i,t},\bar{z}_{i,t})\right\|\leq U,$
@@
$\leq\left\|x_{i,t}-\eta_{i}\nabla_{i}f_{i}(x_{t},\bar{z}_{i,t})-x_{\lambda_{i}}^{*}\right\|^{2}$
$=\left\|x_{i,t}-x_{\lambda_{i}}^{*}\right\|^{2}-2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{i}f_{i}(x_{t},\bar{z}_{i,t})\rangle$
$\quad+\eta_{i}^{2}\left\|\nabla_{i}f_{i}(x_{t},\bar{z}_{i,t})\right\|^{2}$
@@
$\quad+2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{i}f_{i}(x_{t},\bar{z}_{i,t})-\nabla_{i}f_{i}(x_{t},\bar{M}_{i}(x,z_{i,t}))\rangle$
@@
The first inequality stems from the update step of the decisions in Algorithm 1 and the use of fundamental properties of the projection operator. The second equality is obtained by taking into account the property that $\nabla_{x_{i}}\hat{H}_{i}(x_{t},z_{i,t})=\nabla_{i}f_{i}(x_{t},\bar{M}_{i}(x,z_{i,t}))$. The second inequality is obtained by applying the Lipschitz property of $\nabla_{i}f_{i}$, while the last one from the fact that the maximizer of the inner loop in Algorithm 1 calculates the solution from an initial point $z_{i,t}$ with $\epsilon_{i}$ accuracy. Taking expectations on both sides with respect to $z_{i,t}$, we have
@@
$\leq\mathbb{E}\left\|x_{i,t}-x_{\lambda_{i}}^{*}\right\|^{2}-2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{i}H_{i}(x_{t})\rangle+\eta_{i}^{2}U^{2}+$
@@
$\leq\mathbb{E}\left\|x_{i,t}-x_{\lambda_{i}}^{*}\right\|^{2}-2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{i}H_{i}(x_{t})-\nabla_{i}H_{i}(x_{\lambda}^{*})\rangle$
@@
$\leq\mathbb{E}\left\|x_{t}-x_{\lambda}^{*}\right\|^{2}-2\eta\sum_{i=1}^{N}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{i}H_{i}(x_{t})-\nabla_{i}H_{i}(x_{\lambda}^{*})\rangle$

### v2

@@ Abstract
We study a class of distributionally robust games where agents are allowed to heterogeneously choose their risk aversion with respect to distributional shifts of the uncertainty. In our formulation, heterogeneous Wasserstein ball constraints on each distribution are enforced through a penalty function leveraging a Lagrangian formulation. We then formulate the distributionally robust game as a variational inequality problem, and show that under certain assumptions the original seemingly infinite-dimensional Nash equilibrium problem is equivalent to a multi-agent but finite-dimensional variational inequality problem with a strongly monotone mapping. Due to the inner maximization problem, it is however still challenging to calculate a distributionally robust Nash equilibrium. To this end, we design an approximate Nash equilibrium seeking algorithm and prove convergence of the average regret to a quantity that diminishes with the number of iterations, thus learning the desired equilibrium up to an a priori specified accuracy. Numerical simulations corroborate our theoretical findings.
@@ 1.1 Motivation
Modern engineering applications are increasingly operated by autonomous agents that interact strategically to optimize interdependent objectives subject to constraints, thus giving rise to a game-theoretic setting. Recent developments in control, optimization, and machine learning rely heavily on a data-driven multi-agent framework. As some examples, power systems Sun2015; Molzahn2017 and networked robotics Bullo2009, Cortes2017, typically involve a high number of interacting decision-makers whose objectives are coupled through shared resources or performance constraints. At the same time, the environment these agents operate in is often uncertain, and its accurate probabilistic description is rarely known in practice. Thus, probabilistic models are often estimated from a finite collection of data, while statistical and algorithmic analyses are performed for the proposed methods. For uncertain multi-agent games fabiani_2022_automatica; Fele2019; Fele_CDC_2019; Paccagnan_Campi_2019; PantazisFeleMargellos_Automatica2024 perform a detailed distribution-free probabilistic analysis for different game settings. From an algorithmic viewpoint, FranciGrammatico_TAC2021; FranciGrammatico_Automatica2022 design distributed schemes for stochastic generalized Nash equilibrium seeking. However, modeling uncertainty based on an empirical/ estimated distributions calculated on the basis of available samples and treated as ground truth can lead to lack of robustness towards even modest distribution shifts caused, for example, by changing operating conditions, sensor noise, seasonal demand variations or adversarial attacks. Thus, the decisions made are no longer (almost) optimal or as robust as expected.
@@
Distributionally robust optimization (DRO) was developed precisely to tackle this mismatch between the true distribution and its estimated counterpart. In particular, instead of optimizing for one nominal distribution, DRO optimizes for the worst-case distribution within a considered ambiguity set. This equips the decision with quantifiable out-of-sample and finite-sample performance guarantees. Thanks to duality and optimal transport tools EsfahaniKuhn2018; AghassiBertsimas2006 even when the ambiguity set is a Wasserstein ball centered at the empirical distribution, the resulting reformulation can be tractable, oftentimes formulated in the form of large linear programs under relatively mild assumptions. Classical DRO, however, is centralized, i.e., it assumes a single decision-maker and a single source of data. In contrast, in networked and learning-enabled systems, data are heterogeneous and private. Furthermore, as certain agents have different access to information/data, they might exhibit different levels of risk-aversion. A significant challenge that arises in several multi-agent systems is that agents can exhibit self-interested behaviour, wishing to minimize their own cost instead of a common social welfare cost. This framework gives rise to heterogeneous distributionally robust games, where each agent hedges against their own distributional ambiguity while interacting strategically with others.
@@
In wang_icml22, the authors develop a no-regret learning algorithm for multi-agent convex games under CVaR-based risk measures. The work in Wang_tac_2025, studies convex games with heterogeneous feedback, while Wang_acc24 develops learning dynamics that converge to Nash equilibria in risk-averse games, where players optimize CVaR-type objectives. The work PantazisFranciGrammatico2023 relied on epigraphic reformulation methods to provide tractable reformulations for such games. When the agents are affected by a common ambiguity set based on shared data and ambiguity radius, the results in PantazisFranciGrammatico2023 recast the original DRNEP in a form which can provably be solved efficiently via centralized, decentralized or distributed algorithmic schemes. However, when ambiguity sets are private, this approach often leads to constraints which are private and coupling. An approach to fix this issue is to share the private constraints among agents. Unfortunately, this would introduce further cooperation among players, thus changing the original non-cooperative game with no clear connection between the equilibria of the two problems. To solve this problem, PantazisFranciGrammatico2023 propose a primal-dual algorithm inspired by the golden-ratio method, malitsky2020golden, Franci2022_goldenratio, to empirically converge to an equilibrium of the original problem. Following a different approach, Pantazis_dr_games_2024 focus on a particular class of DRNEPs and recast the problem into a variational inequality and propose efficient algorithmic methods to calculate an equilibrium. The recent paper in Parsi2025 follows again a VI formulation, but instead proposes a gradient descent/ascent for which regret guarantees can be obtained under certain assumptions. lanzetti2025 introduce an alternative equilibrium concept based on optimal transport that protects against uncertainty in opponents’ behavior, while fabiani_franci_2023 study distributionally robust generalized Nash games over Wasserstein distributionally robust chance-constraints.
@@ 1.2 Main contributions
    Lagrangian formulation of DRNEPs with heterogeneous ambiguity sets: To improve the computational efficiency in DRNE equilibrium seeking, while still maintaining certain distributional robustness properties, we propose a Lagrangian formulation of DRNEPs, where each agent enforces its Wasserstein ambiguity constraint via a penalty term in their cost, thus extending the formulation of the single-agent setting proposed in sinha2017certifying. Our formulation enables heterogeneous and private ambiguity sets and risk-aversion levels without sharing constraints or data, while preserving the noncooperative nature of the game. We formalize the distributionally robust Nash equilibria (DRNE) for this novel setting and establish a connection with the DRNE of the original distributionally robust game. In our setting the agents’ penalty parameters act as individual distributional robustness-tuning coefficients.
@@ 2 Problem Formulation
Note that the inner supremum in Definition 2 involves selecting an adversarial probability distribution that maximizes the cost of each agent. The seemingly infinite dimensional problem $(G)$ can be shown to admit tractable reformulations through duality and the use of epigraphic reformulations in PantazisFranciGrammatico2023 based on a methodology in the same spirit of EsfahaniKuhn2018. In particular, leveraging Kantorovich–Rubinstein duality, $(G)$ admits under mild assumptions the following reformulation:
@@
$\min_{x_{i}\in X_{i}}\max_{\mathbb{P}_{i}}\mathbb{E}_{\xi_{i}\sim\mathbb{P}_{i}}[f_{i}(x_{i},x_{-i},\xi_{i})]-\lambda_{i}W_{p}(\hat{\mathbb{P}}_{K_{i}},\mathbb{P}_{i})$ (6)
@@
Proof: The proof follows by adaptation of sinha2017certifying for each agent $i\in\mathcal{N}$, given $x_{-i}$. $\blacksquare$
@@
For the game $G_{L}$ in (6) with $\lambda=\operatorname*{col}(\lambda_{i})_{i=1}^{N}$, a point $x^{\ast}_{\lambda}$ is a distributionally robust Nash equilibrium (DRNE) if
$H_{i}(x_{\lambda}^{\ast})$ $\leq\min_{x_{i}\in X_{i}}H_{i}(x_{i},x^{\ast}_{-i,\lambda}),$ (8)
@@ 3.1 Distributionally Robust Nash Equilbrium Problems as a Variational Inequality Problem
In this section we recast the game $G_{L}$ in (6) as a variational inequality and establish the conditions that guarantee (strong) monotonicity of the corresponding mapping. We propose an equilibrium seeking algorithm to calculate efficiently a DRNE of (7). In particular, we propose Algorithm 1 which is based on an inexact proximal operation to calculate the inner supremum followed by a projected gradient descent for each agent $i\in\mathcal{N}$. We show that under certain conditions, convergence of the iterations to a neighbourhood of the equilibrium is guaranteed.
@@
Under Assumption 3.1a), $h_{i}(x,\xi_{i},\hat{\xi}_{i})$ is $2\lambda_{i}$-strongly concave in $\xi_{i}$. $\square$
@@
Given Assumptions  3.1 and 3.1, $H_{i}$ is differentiable and satisfies
@@
Proof: The proof is an adaptation of Lemma 1 in sinha2017certifying applied to the cost function of each agent $i\in\mathcal{N}$. $\blacksquare$
@@
[…]
Proof: The proof can be obtained by applying Proposition 1.4.2 in FachineiPang to VI($F_{K},X$). $\blacksquare$
@@
where $\xi=\operatorname*{col}((\xi_{i})_{i=1}^{N})$. We then consider the following assumption:
@@
$G$ is $\mu$-strongly monotone in its first argument for any $\xi_{i}\in\Xi_{i}$, $i\in\mathcal{N}$, i.e, for any decisions $x,y\in X$ and $\xi_{i}\in\Xi_{i}$ it holds that
@@
$\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{x_{i}}H_{i}(x,\hat{\xi}_{i})-\nabla_{x_{i}}H_{i}(y,\hat{\xi}_{i}),x_{i}-y_{i}\Big\rangle$
$=\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{x_{i}}f_{i}(x,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i})),x_{i}-y_{i}\Big\rangle$
$=\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{x_{i}}f_{i}(x,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))$
$\quad+\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i})),x_{i}-y_{i}\Big\rangle$
@@
$+\sum_{i=1}^{n}\mathbb{E}_{\hat{\xi}_{i}\sim\hat{\mathbb{P}}_{i}}\Big\langle\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i})),x_{i}-y_{i}\Big\rangle$
@@
$\left\|\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(x,\hat{\xi}_{i}))-\nabla_{x_{i}}f_{i}(y,\bar{M}_{i}(y,\hat{\xi}_{i}))\right\|$
@@ 3.2 Average Convergence Guarantees
Let Assumptions 3.1 and 3.1, 3.1 hold. Then, the iterates $(x_{t})_{t\in\mathbb{N}}$ of Algorithm 1 averagely converge to an $\epsilon$-DRNE, i.e.,
@@
$\left\|\nabla_{x_{i}}f_{i}(x_{i,t},x_{-i,t},\bar{z}_{i,t})\right\|\leq U,$
@@
$\leq\left\|x_{i,t}-\eta_{i}\nabla_{x_{i}}f_{i}(x_{t},\bar{z}_{i,t})-x_{\lambda_{i}}^{*}\right\|^{2}$
$=\left\|x_{i,t}-x_{\lambda_{i}}^{*}\right\|^{2}-2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{x_{i}}f_{i}(x_{t},\bar{z}_{i,t})\rangle$
$\quad+\eta_{i}^{2}\left\|\nabla_{x_{i}}f_{i}(x_{t},\bar{z}_{i,t})\right\|^{2}$
@@
$\quad+2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{x_{i}}f_{i}(x_{t},\bar{z}_{i,t})-\nabla_{x_{i}}f_{i}(x_{t},\bar{M}_{i}(x,z_{i,t}))\rangle$
@@
The first inequality stems from the update step of the decisions in Algorithm 1 and the use of fundamental properties of the projection operator. The second equality is obtained by taking into account the property that
$\nabla_{x_{i}}\hat{H}_{i}(x_{t},z_{i,t})=\nabla_{x_{i}}f_{i}(x_{t},\bar{M}_{i}(x,z_{i,t})).$
The second inequality is obtained by applying the Lipschitz property of $\nabla_{x_{i}}f_{i}$, while the last one from the fact that the maximizer of the inner loop in Algorithm 1 calculates the solution from an initial point $z_{i,t}$ with $\epsilon_{i}$ accuracy. Taking expectations on both sides with respect to $z_{i,t}$, we have
@@
$\leq\mathbb{E}\left\|x_{i,t}-x_{\lambda_{i}}^{*}\right\|^{2}-2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{x_{i}}H_{i}(x_{t})\rangle+\eta_{i}^{2}U^{2}+$
@@
$\leq\mathbb{E}\left\|x_{i,t}-x_{\lambda_{i}}^{*}\right\|^{2}-2\eta_{i}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{x_{i}}H_{i}(x_{t})-\nabla_{x_{i}}H_{i}(x_{\lambda}^{*})\rangle$
@@
$\leq\mathbb{E}\left\|x_{t}-x_{\lambda}^{*}\right\|^{2}-2\eta\sum_{i=1}^{N}\langle x_{i,t}-x_{\lambda_{i}}^{*},\nabla_{x_{i}}H_{i}(x_{t})-\nabla_{x_{i}}H_{i}(x_{\lambda}^{*})\rangle$